Can someone please help me with working out this problem: 8x^2-10x+3
I can not figure it out with there being no common factor. I need to factor the problem completely. Thank you in advance!

Answer by checkley71(8403)

8X~2-10X+3=0 OR (2X-1)(4X-3)=0 OR 2X-1=0 OR 2X=1 OR X=.5 & 4X-3=0 OR 4X=3 OR X=3/4 OR X=.75